Construct a confidence interval of the population proportion at the given level of confidence. x = 120, n = 1200, 95% confide and i need help finding the upper bound confidence interval as well
Answer #1

Solution :

Given that x = 120, n = 1200

=> p = x/n

= 120/1200

= 0.1

=> q = 1 - p = 0.9

=> For 95% confidence level, Z = 1.96

=> The lower bound of the confidence interval is

=> p - Z*sqrt(p*q/n)

=> 0.1 - 1.96*sqrt(0.1*0.9/120)

=> 0.0463

=> 0.046 (rounded)

=> The upper bound of the confidence interval is

=> p + Z*sqrt(p*q/n)

=> 0.1 + 1.96*sqrt(0.1*0.9/120)

=> 0.1537

=> 0.154 (rounded)